OpenStudy (andu1854):

(cos^3x - sin^3x)/(cosx - sinx) = 1 + sinx cosx; Verify the Identity. I am not sure how to get to the answer. Any help getting me started would be greatly appreciated.

OpenStudy (tkhunny):

You should recognize the form of the numerator. A difference of cubes. This can be factored.

OpenStudy (andu1854):

Can you factor out a sincos to then create sincos(sin^2 +cos^2

OpenStudy (anonymous):

\(a^3-b^3=(a-b)(a^2+ab+b^2)\)

OpenStudy (anonymous):

As tkhunny mentioned, you have something that looks like \[\frac{a^3-b^3}{a-b}\] Using the difference of cubes pattern, you can factor to get \[\frac{(a-b)(a^2+ab+b^2)}{(a-b)}\]

OpenStudy (andu1854):

Thank you very much... let me see what I do next, but that was very cool...

OpenStudy (andu1854):

Thank you for the help on this one... I got it...